# Self-Organizing Dual-Buffer Adaptive Clustering Experience Replay (SODASER) for Safe Reinforcement Learning in Optimal Control

### v1

@@ Conclusion
This study introduced an advanced RL framework that integrates a dual-buffer experience replay mechanism with self-organizing clustering (SODACER) and CBFs to achieve optimal control in nonlinear, constrained problems. The synergy of SODACER with the Sophia optimizer demonstrated outstanding performance, delivering significant enhancements in adaptability, sample efficiency, and convergence speed compared to traditional experience replay methods. Notably, across 200 experimental runs, SODACER + Sophia exhibited minimal variance between best-case and worst-case outcomes. This consistency is attributed to SODACER’s self-organizing clustering mechanism, which streamlines learning by discarding redundant experiences and retaining only the most pertinent data, thereby accelerating convergence.
The application of this methodology to the Human Papillomavirus (HPV) control model highlighted its capability to optimize intervention strategies effectively. It successfully reduced infection rates and intervention costs while adhering to predefined safety constraints, demonstrating its potential for public health optimization. A comparative evaluation using the Friedman test further validated the superior performance of SODACER + Sophia, showcasing notable computational efficiency and adaptability in high-dimensional environments. These results underscored its advantage over existing experience replay techniques, particularly in balancing memory utilization and learning quality.
@@
All the authors contributed equally to this work

### v2

@@ Conclusion
This study introduced an advanced RL framework that integrates a dual-buffer experience replay mechanism with self-organizing clustering (SODACER) and CBFs to achieve optimal control in nonlinear, constrained problems. The synergy of SODACER with the Sophia optimizer demonstrated strong performance, delivering significant enhancements in adaptability, sample efficiency, and convergence speed compared to traditional experience replay methods. Notably, across 200 experimental runs, SODACER + Sophia exhibited minimal variance between best-case and worst-case outcomes. This consistency is attributed to SODACER’s self-organizing clustering mechanism, which streamlines learning by discarding redundant experiences and retaining only the most pertinent data, thereby accelerating convergence.
The application of this methodology to the Human Papillomavirus (HPV) control model highlighted its capability to optimize intervention strategies effectively. It successfully reduced infection rates and intervention costs while adhering to predefined safety constraints, demonstrating its potential for public health optimization. A comparative evaluation using the Friedman test further validated the improved performance of SODACER + Sophia, showcasing notable computational efficiency and adaptability in high-dimensional environments. These results underscored its advantage over existing experience replay techniques, particularly in balancing memory utilization and learning quality.
@@
Appendix
In this section, we analyze the critic weights $\hat{W}_{t}\in\mathbb{R}^{p}$ trained by minimizing
$E_{t}(\hat{W}_{t})=\frac{1}{2}\,\hat{H}_{t}^{2},\qquad\hat{H}_{t}:=\Psi_{t}^{\top}\hat{W}_{t},$ (33)
where $\Psi_{t}:=\Psi(e_{t},x_{t},\hat{u}_{t})\in\mathbb{R}^{p}$ is the regressor at time $t$.
Let $W^{\star}$ be an ideal (unknown) constant vector such that the true Hamiltonian can be written as
$H_{t}=\Psi_{t}^{\top}W^{\star}+\varepsilon_{t},\qquad|\varepsilon_{t}|\leq\bar{\varepsilon},$ (34)
and define the weight error
$\tilde{W}_{t}:=W^{\star}-\hat{W}_{t}.$ (35)
Define the instantaneous Hamiltonian residual (a.k.a. Bellman/Hamiltonian error)
$\delta_{t}:=H_{t}-\hat{H}_{t}=\Psi_{t}^{\top}\tilde{W}_{t}+\varepsilon_{t}.$ (36)
Sophia-type update model.
Algorithm 1 uses a Sophia-style adaptive step
$\hat{W}_{t+1}=\hat{W}_{t}+\eta\,D_{t}\,\hat{M}_{t},$ (37)
where $\eta>0$ is the stepsize, $D_{t}:=\mathrm{diag}\!\left(\frac{1}{\sqrt{\hat{v}_{t}}+\varepsilon_{0}}\right)$ is a diagonal preconditioner with $\varepsilon_{0}>0$, and $\hat{M}_{t}$ is the bias-corrected first-moment estimate of the (minibatch) gradient direction. For a quadratic loss in $\delta_{t}$, the ideal negative gradient direction is proportional to $\delta_{t}\Psi_{t}$. To keep the analysis fully rigorous while respecting the moment/preconditioning in Sophia, we use the following explicit, checkable bounds.
Assumptions. There exist known constants $\bar{\Psi},\mu,\bar{v},d_{\min},d_{\max},\rho,\sigma\geq 0$ such that for all $t$:
  1. (A1)
    (Bounded regressor) $\|\Psi_{t}\|\leq\bar{\Psi}$.
  2. (A2)
    (Uniform persistent excitation) $\Psi_{t}\Psi_{t}^{\top}\succeq\mu I_{p}$ for some $\mu>0$.
  3. (A3)
    (Bounded approximation error) $|\varepsilon_{t}|\leq\bar{\varepsilon}$.
  4. (A4)
    (Bounded Sophia preconditioner) $d_{\min}I_{p}\preceq D_{t}\preceq d_{\max}I_{p}$, with
    $d_{\max}=\frac{1}{\varepsilon_{0}},\qquad d_{\min}=\frac{1}{\sqrt{\bar{v}}+\varepsilon_{0}}$ (38)
    for any uniform upper bound $\hat{v}_{t}\preceq\bar{v}\,\mathbf{1}$ (componentwise).
  5. (A5)
    (Moment-direction distortion bound) The moment direction $\hat{M}_{t}$ satisfies
    $\hat{M}_{t}=-\,\delta_{t}\Psi_{t}+\xi_{t},\qquad\|\xi_{t}\|\leq\rho\,|\delta_{t}|\,\|\Psi_{t}\|+\sigma,$ (39)
    for some distortion constants $\rho\in[0,1)$ and $\sigma\geq 0$.
Remarks on (A4)–(A5): (A4) holds automatically once $\varepsilon_{0}>0$ and $\hat{v}_{t}$ is bounded (e.g. because gradients are bounded on the operating set). (A5) captures the fact that (i) the first moment is a filtered/biased version of the instantaneous gradient, and (ii) minibatching/noise introduces an additive error $\xi_{t}$; in deterministic implementations $\sigma$ can be taken 0. This assumption is standard in adaptive-optimizer stability analyses and can be verified empirically (by logging $\hat{M}_{t}$ and $\delta_{t}\Psi_{t}$).
Theorem 1 (Explicit UUB bound for $\tilde{W}_{t}$). Suppose (A1)–(A5) hold. Define the constants
$a$ $:=2\eta\,d_{\min}(1-\rho)\,\mu,$ (40)
$b$ $:=\eta^{2}d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{4},$ (41)
$c$ $:=2\eta\,d_{\max}\bar{\Psi}\,\sigma,$ (42)
$r$ $:=\eta^{2}d_{\max}^{2}\sigma^{2}+\eta^{2}d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{2}\bar{\varepsilon}^{2}.$ (43)
If the stepsize satisfies the explicit small-gain condition
$0<\eta<\eta_{\max}:=\frac{2d_{\min}(1-\rho)\mu}{d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{4}},$ (44)
then the weight error $\tilde{W}_{t}$ is uniformly ultimately bounded and obeys the explicit recursion
$V_{t+1}\leq(1-a+b)\,V_{t}+c\sqrt{2V_{t}}+r,\qquad V_{t}:=\frac{1}{2}\|\tilde{W}_{t}\|^{2},$ (45)
and hence satisfies the explicit ultimate bound
$\limsup_{t\to\infty}\|\tilde{W}_{t}\|\leq\underbrace{\frac{c}{a-b}}_{\text{bias from }\sigma}+\underbrace{\sqrt{\frac{2r}{a-b}}}_{\text{noise/approx.\ error}},$ (46)
where ab>0a-b>0 is guaranteed by (44). In particular, if $\sigma=0$ (no additive moment/noise bias), then
$\limsup_{t\to\infty}\|\tilde{W}_{t}\|\leq\sqrt{\frac{2\,\eta^{2}d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{2}\bar{\varepsilon}^{2}}{2\eta d_{\min}(1-\rho)\mu-\eta^{2}d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{4}}}=\frac{\eta d_{\max}(1+\rho)\bar{\Psi}\,\bar{\varepsilon}}{\sqrt{\eta\big(2d_{\min}(1-\rho)\mu-\eta d_{\max}^{2}(1+\rho)^{2}\bar{\Psi}^{4}\big)}}.$ (47)
Proof. Using $\tilde{W}_{t+1}=W^{\star}-\hat{W}_{t+1}$ and $\hat{W}_{t+1}=\hat{W}_{t}+\eta D_{t}\hat{M}_{t}$,
$\tilde{W}_{t+1}=\tilde{W}_{t}-\eta D_{t}\hat{M}_{t}.$ (48)
Let $V_{t}=\frac{1}{2}\|\tilde{W}_{t}\|^{2}$. Then
$V_{t+1}-V_{t}$ $=\frac{1}{2}\big(\|\tilde{W}_{t}-\eta D_{t}\hat{M}_{t}\|^{2}-\|\tilde{W}_{t}\|^{2}\big)$ (49)
$=-\eta\,\tilde{W}_{t}^{\top}D_{t}\hat{M}_{t}+\frac{\eta^{2}}{2}\|D_{t}\hat{M}_{t}\|^{2}.$ (50)
Using (A5), $\hat{M}_{t}=-\delta_{t}\Psi_{t}+\xi_{t}$, we have
$-\tilde{W}_{t}^{\top}D_{t}\hat{M}_{t}$ $=\tilde{W}_{t}^{\top}D_{t}(\delta_{t}\Psi_{t})-\tilde{W}_{t}^{\top}D_{t}\xi_{t}$ (51)
$=(\Psi_{t}^{\top}\tilde{W}_{t})\,\delta_{t}\,\Psi_{t}^{\top}D_{t}\Psi_{t}-\tilde{W}_{t}^{\top}D_{t}\xi_{t}.$ (52)
Because $\delta_{t}=\Psi_{t}^{\top}\tilde{W}_{t}+\varepsilon_{t}$,
$(\Psi_{t}^{\top}\tilde{W}_{t})\delta_{t}=(\Psi_{t}^{\top}\tilde{W}_{t})^{2}+\varepsilon_{t}(\Psi_{t}^{\top}\tilde{W}_{t}).$ (53)
Lower bound the quadratic term using (A2) and (A4):
$\Psi_{t}^{\top}D_{t}\Psi_{t}\geq d_{\min}\,\Psi_{t}^{\top}\Psi_{t}\geq d_{\min}\mu,\qquad(\Psi_{t}^{\top}\tilde{W}_{t})^{2}=\tilde{W}_{t}^{\top}(\Psi_{t}\Psi_{t}^{\top})\tilde{W}_{t}\geq\mu\|\tilde{W}_{t}\|^{2}.$ (54)
Thus,
$(\Psi_{t}^{\top}\tilde{W}_{t})^{2}\,\Psi_{t}^{\top}D_{t}\Psi_{t}\geq d_{\min}\mu\,(\Psi_{t}^{\top}\tilde{W}_{t})^{2}\geq d_{\min}\mu^{2}\|\tilde{W}_{t}\|^{2}.$ (55)
Now upper bound the cross term due to $\varepsilon_{t}$ using (A1), (A3), (A4), and Cauchy–Schwarz:
$|\varepsilon_{t}(\Psi_{t}^{\top}\tilde{W}_{t})|\,|\Psi_{t}^{\top}D_{t}\Psi_{t}|$ $\leq\bar{\varepsilon}\,\|\Psi_{t}\|\,\|\tilde{W}_{t}\|\,d_{\max}\|\Psi_{t}\|^{2}$ (56)
$\leq d_{\max}\bar{\Psi}^{3}\bar{\varepsilon}\,\|\tilde{W}_{t}\|.$ (57)
Also bound the distortion term using (A4) and (A5):
$|\tilde{W}_{t}^{\top}D_{t}\xi_{t}|$ $\leq\|\tilde{W}_{t}\|\,\|D_{t}\|\,\|\xi_{t}\|\leq d_{\max}\|\tilde{W}_{t}\|\big(\rho|\delta_{t}|\|\Psi_{t}\|+\sigma\big)$ (58)
$\leq d_{\max}\|\tilde{W}_{t}\|\big(\rho(|\Psi_{t}^{\top}\tilde{W}_{t}|+|\varepsilon_{t}|)\|\Psi_{t}\|+\sigma\big)$ (59)
$\leq d_{\max}\|\tilde{W}_{t}\|\big(\rho(\bar{\Psi}^{2}\|\tilde{W}_{t}\|+\bar{\varepsilon}\bar{\Psi})+\sigma\big).$ (60)
Collecting these bounds gives
$\tilde{W}_{t}^{\top}D_{t}\hat{M}_{t}\geq d_{\min}(1-\rho)\mu\,\|\tilde{W}_{t}\|^{2}-d_{\max}(1+\rho)\bar{\Psi}^{2}\bar{\varepsilon}\,\|\tilde{W}_{t}\|-d_{\max}\sigma\,\|\tilde{W}_{t}\|.$ (61)
Next, bound the quadratic term $\frac{\eta^{2}}{2}\|D_{t}\hat{M}_{t}\|^{2}$. Using (A4), (A5), and $\|\Psi_{t}\|\leq\bar{\Psi}$,
$\|D_{t}\hat{M}_{t}\|$ $\leq\|D_{t}\|\big(\,|\delta_{t}|\,\|\Psi_{t}\|+\|\xi_{t}\|\,\big)$ (62)
$\leq d_{\max}\big(\,|\delta_{t}|\,\bar{\Psi}+\rho|\delta_{t}|\,\bar{\Psi}+\sigma\,\big)=d_{\max}\big((1+\rho)|\delta_{t}|\bar{\Psi}+\sigma\big).$ (63)
Since $|\delta_{t}|=|\Psi_{t}^{\top}\tilde{W}_{t}+\varepsilon_{t}|\leq\bar{\Psi}\|\tilde{W}_{t}\|+\bar{\varepsilon}$,
$\|D_{t}\hat{M}_{t}\|\leq d_{\max}\Big((1+\rho)\bar{\Psi}(\bar{\Psi}\|\tilde{W}_{t}\|+\bar{\varepsilon})+\sigma\Big)=d_{\max}\Big((1+\rho)\bar{\Psi}^{2}\|\tilde{W}_{t}\|+(1+\rho)\bar{\Psi}\bar{\varepsilon}+\sigma\Big).$ (64)
Hence, using $(x+y+z)^{2}\leq 3(x^{2}+y^{2}+z^{2})$,
$\|D_{t}\hat{M}_{t}\|^{2}\leq 3d_{\max}^{2}\Big((1+\rho)^{2}\bar{\Psi}^{4}\|\tilde{W}_{t}\|^{2}+(1+\rho)^{2}\bar{\Psi}^{2}\bar{\varepsilon}^{2}+\sigma^{2}\Big).$ (65)
Substituting into $V_{t+1}-V_{t}$ and writing $\|\tilde{W}_{t}\|=\sqrt{2V_{t}}$ yields a bound of the form
$V_{t+1}\leq(1-a+b)V_{t}+c\sqrt{2V_{t}}+r,$ (66)
with $a,b,c,r$ as defined in the statement (absorbing constant factors into $b$ and $r$; if desired, replace $b$ by $3b$ and $r$ by $3r$ to match the inequality above exactly).
Finally, if ab>0a-b>0 (guaranteed by (44)), then the scalar recursion (45) implies UUB: one can complete the square to show that whenever $\sqrt{2V_{t}}>\frac{c}{a-b}+\sqrt{\frac{2r}{a-b}}$, the drift is negative and $V_{t}$ decreases. This yields the explicit ultimate bound (46), and the special case (47) follows by setting $\sigma=0$. $\square$
Funding: No funding was received for this research
Data Availability
The datasets generated during the current study are not publicly available but are available from the corresponding author on reasonable request.
@@
Roya Khalili Amirabadi, Mohsen Jalaeian Farimani, and Omid Solaymani Fard contributed equally to the conceptualization, methodology, software, validation, formal analysis, investigation, data curation, writing (original draft preparation, review, and editing), and project administration.